Could you please help me out?
The perimeter of a room is 96 feet. The room is rectangular in shape and the length of the room is 8 feet more than the width. What is the area of the room?

Answer by checkley77(12569)

2X+2(X+4)=96
2X+2X+8=96
4X=96-8
4X=88
X=88/4
X=22 THE WIDTH OF THE ROOM.
22=4=26 FOR THE LENGTH OF THE ROOM.
PROOF:
2*22=2*26=96
44+52=96
96=96
AREA=22*26=572 FEET^2 IS THE AREA OF THE ROOM.